# A Nonparametric Bayesian Solution of the Empirical Stochastic Inverse Problem

### v1

@@ 1.5 Outline of the paper
In Sec. 2, we develop the formulation of the Bayesian solution of the SIP using disintegration of measures. We describe conditions that guarantee the solution can be expressed in terms of a conditional density in Sec. 2.4. We describe the maximum entropy property of the solution corresponding to the uniform prior in Sec. 2.5. In Sec. 2.6, we describe the continuity property of the solution. In Sec. 3, we describe and analyze the numerical solution method for the eSIP based on reweighting of random samples. In Sec. 4, we present two examples. We develop an alternative accept-reject numerical solution method in Sec. 5. We present a conclusion in Sec. 6. Proofs and additional theoretical developments are provided in Appendix A.
@@ 2 Solution of the Stochastic Inverse Problem
In developing the Bayesian solution of the SIP, we move from the most general formulation of the SIP to a practical formulation by layering the assumptions required for the development. The development is necessarily abstract because the intent is to establish the theoretical framework that provides the basis to compute a robustly accurate empirical approximation of a posterior distribution solving the eSIP that is conditioned on experimental data and the choice of prior.
@@ 2.1 Disintegration of measures
Figure 4: Left: we illustrate disintegration when the conditional probabilities are written in terms of densities. The probability on $\Lambda$ is represented by a density shaded in green and the probability on $\mathcal{D}$ is represented by a density shaded in blue. The disintegrated density on $Q^{-1}(q)$ is indicated with a darker shade of green. Right: we illustrate the numerical solution method by random sampling. The observed data is indicated by dark red points on $\mathcal{D}$ and the associated empirical distribution is shown in blue relative to the partition $\{I_{i}\}$. We show the samples $\{\lambda_{i}\}$ in $\Lambda$ as points and $Q^{-1}(I_{i})$ is shaded in gold. Figure is adapted from [46].
@@ 2.2 Bayesian solution of the SIP
To obtain observational data for the eSIP, we compute 200 draws from a Bernoulli random variable with probability $\frac{61}{90}$ of 11 and use these to compute the estimate $\begin{pmatrix}\hat{P}_{\mathcal{D}}(0)\\ \hat{P}_{\mathcal{D}}(1)\end{pmatrix}=\begin{pmatrix}.34\\ .66\end{pmatrix}$. We compute the solution of the eSIP corresponding to $\hat{P}_{\mathcal{D}}$. For reasons explained below, in the absence of prior information, we assume a uniform prior,
@@
where $A$ is any set in $\Lambda$ and $\chi_{A}$ is the indicator function for set $A$, as well as the estimated solution of the eSIP,
@@
In order to express the disintegration (2.5) in concrete terms, we develop the properties of $Q^{-1}$. In addition to Assumption 2.1, we assume the following holds.
@@
The compactness of $\Lambda$ is reasonable because parameters are generally bounded for most process models [^2] We emphasize that $\Lambda$ is not the domain for parameters in a parameterized family of probability distributions.. In fact, determining $\Lambda$ is often an important part of constructing a model, see [46] where the choice of $\Lambda$ for transmission rates for different strains of COVID-19 is discussed. Assuming $\overline{\mathrm{int}\,(\Lambda)}=\Lambda$ and $\mu_{\mathcal{L}}(\partial\Lambda)=0$ are technical assumptions required for approximate solution by finite sampling. We illustrate in Fig. 5. The smoothness assumption on $Q$ can be established for broad classes of models since it is typically assumed for numerical solution of the model. The assumption of GD means that $\Lambda$ can be covered by a finite number of open sets up to a set of measure 0 such that $J_{Q}$ has full rank $m$ on each set. In these sets, no component of $J_{Q}$ can be expressed as a linear combination of its other components. If this is violated, $Q$ should be simplified to obtain a new $Q$ map into a range with smaller dimension.
@@
2.4 Writing the solution of the SIP in terms of a conditional density
We develop a formula for the solution in terms of a conditional density function. We assume,
@@ 2.4 Writing the solution of the SIP in terms of a conditional density
Assume Assumptions 2.1, 2.5, 2.7, and 2.8 hold. There is a solution of the SIP ${P}_{\Lambda}$ that is absolutely continuous with respect to $\mu_{\Lambda}$ and $dP_{\Lambda}={\rho}_{\Lambda}\,d\mu_{\Lambda}$ with
@@
2.5 The maximum entropy property of the uniform prior solution
The solution of the SIP corresponding to the uniform prior $P_{p}=\frac{1}{\mu_{\Lambda}(\Lambda)}\mu_{\Lambda}$ has density,
@@ 2.5 The maximum entropy property of the uniform prior solution
We call the solution corresponding to the uniform prior the posterior solution conditioned on the uniform prior or the uniform prior (posterior) solution.
The uniform prior has a special property that makes it the default choice for solving the SIP in the absence of an informed prior. Namely, it is an expression of the Principle of Insufficient Reason. The translation invariance of the uniform prior respects the fact that since distinct points on the generalized contour $Q^{-1}(q)$ cannot be distinguished using data on $Q$, the probability of points on $Q^{-1}(q)$ should be equal. The next theorem shows that the solution ${P}_{\Lambda}$ corresponding to the uniform prior is the least biased solution of the SIP in a certain sense. We give the proof in § A.3.
@@
Assume Assumptions 2.1, 2.5, and 2.7 hold. The solution of the SIP corresponding to the uniform prior has maximum entropy among all solutions that are absolutely continuous with respect to $\mu_{\Lambda}$ and whose prior is absolutely continuous with respect to the Lebesgue measure.
@@
In Example 3, the entropy of the SIP solution for the uniform prior is 2.1746 and the entropy for the solution for the prior (2.7) is 1.9805.
2.6 Continuity
We develop sufficient conditions that guarantee the density of the solution of the SIP is continuous a.e. A key ingredient is the a.e. continuity of the surface (Hausdorff) measure of the generalized contours. This depends on the interaction of the generalized contours with the boundary of $\Lambda$. For example, if the boundary has measure larger than 0, there is little chance that the surface measure of the contours is continuous, see Fig. 5.
@@ 2.6 Continuity
To analyze the interaction of the generalized contours with the boundary of $\Lambda$, we assume the boundary is a piecewise smooth manifold. Specifically, we assume
@@
Assume Assumptions 2.1, 2.5, 2.7, 2.8, and 2.11 hold and the density of the prior $\rho_{\mathrm{p}}$ is continuous a.e. Then, $\tilde{\rho}_{\mathrm{p},\mathcal{D}}(q)$ is continuous in a neighborhood of each point $q\in\mathcal{D}$. In addition, if the density function $\rho_{\mathcal{D}}$ is continuous a.e., then the density ${\rho}_{\Lambda}$ of the solution $P_{\Lambda}$ corresponding to the prior $P_{\mathrm{p}}$ is continuous in a neighborhood of each point $q\in\mathcal{D}$.
@@
3 Numerical solution of the eSIP by reweighting random samples
In this section, we present and analyze a numerical solution method for the eSIP based on random sampling. The convergence analysis is conducted under the assumptions in Theorem 2.9 and the a.e. continuity of the solution is crucial.
@@ 3.2 Constructing the estimate
There are three discretizations involved with the eSIP; the number of observed data points $K$, the number of cells $M$ in a partition of $\mathcal{D}$, and the number of cells $N$ in a partition of $\Lambda$. However, limits on the accuracy of empirical density approximations means that $M$ must be related to $K$. We denote the observed data by $\{q_{i}\}_{i=1}^{\infty}$, which we assume is a random sample generated by a Poisson point process with distribution $P_{\mathcal{D}}$. For $K\geq 1$, we construct an approximation using $\{q_{i}\}_{i=1}^{K}$ that is associated with a partition $\mathcal{I}_{M_{K}}$ of $\mathcal{D}$, where the number of cells $M_{K}$ in $\mathcal{I}_{M_{K}}$ is a function of $K$ that tends monotonically to $\infty$ as $K\to\infty$. We assume the family $\{\mathcal{I}_{M_{K}}\}_{K=1}^{\infty}$ satisfies Assumption 3.1. The approximation is also associated with partitions $\{\mathcal{T}_{N_{\jmath}}\}_{\jmath=1}^{\infty}$ of $\Lambda$, which we assume satisfies Assumption 3.1.
@@
The approximation for ${P}_{\Lambda}$ is
@@
We write the method in a form more conducive to analysis,
@@
An illustration of the method is presented in Fig. 4. Beginning with the observed data, we partition the range $\mathcal{D}$ into a collection of cells and use the data to approximate the probability $P_{\mathcal{D}}$ of each cell. The inverse image under $Q$ of each cell in the partition of $\mathcal{D}$ is a “fat” contour. As the cells in the partition of $\mathcal{D}$ become smaller, the fat contours become “thinner”. We compute a sample of points in $\Lambda$ distributed according to the prior. We estimate the probability of the intersection of a fixed set $A$ with a fat contour by taking the ratio of the number of samples in the intersection of a fixed set $A$ with the fat contour and the total number of samples in the fat contour. Finally, we sum up all the estimates over fat contours that intersect $A$, reweighting the estimate in each fat contour with the probability of its corresponding cell in $\mathcal{D}$. The key to convergence of this approach is the a.e. continuity of the surface measure of the contours. When a fat contour is sufficiently thin, the surface measures of almost all the contours in it are approximately the same.
3.3 The basic statistical properties of the estimate
These properties quantify how the estimate converges to the solution of the eSIP determined by the prior as all the discretizations are refined. We use the notation,
@@ 3.3 The basic statistical properties of the estimate
    The estimate is asymptotically unbiased. For $A\in\mathcal{B}_{\Lambda}$ with $\mu_{\Lambda}(\partial A)=0$,
    $\mathbb{E}\big(\hat{P}_{\Lambda,K,N_{\jmath}}(A)\big)=\sum_{i=1}^{M_{K}}\frac{\breve{p}_{i}^{(K)}}{p_{i}^{(K)}}\big(1-(1-p_{i}^{(K)})^{N_{\jmath}})\,P_{\mathcal{D}}\big(I_{i}^{(K)}\big),$ (3.6)
    and $\lim_{K\to\infty}\lim_{\jmath\to\infty}{\mathbb{E}\big(\hat{P}_{\Lambda,K,N_{\jmath}}(A)\big))}={P}_{\Lambda}(A)$.
@@
    The estimate is strongly consistent. For $A\in\mathcal{B}_{\Lambda}$ with $\mu_{\Lambda}(\partial A)=0$,
@@
3.4 Asymptotic properties of the error of the estimate
@@
Assume Assumptions 2.1, 2.5, 2.7, 2.8, 2.11, and 3.1 hold and assume the densities $\rho_{\mathrm{p}}$ and $\rho_{\mathcal{D}}$ are continuous a.e. Then, $\lim_{\jmath\to\infty}\mathbb{E}(T_{3,i,\jmath})=0$ and $\lim_{\jmath\to\infty}\mathrm{Var}\,(T_{3,i,\jmath})=0$.
@@ 3.5 Using other estimators of $\rho_{\mathcal{D}}$
We consider a general estimate $\hat{\rho}_{k,i}=\hat{P}_{\mathcal{D}}\big(I_{i}^{(k)}\big)$ in (3.3), where $\hat{P}_{\mathcal{D}}$ has density $\hat{\rho}_{\mathcal{D}}$ with respect to $\mu_{\mathcal{D}}$. We prove the following theorem in § A.7.
@@ 4 Examples
We present a couple of simple examples that illustrate the results above. Applications of the eSIP to more complex models with higher dimensional parameter spaces can be found in [10, 12, 14, 13, 52, 35, 15, 16, 18, 53, 6, 42, 46]. The numerical solutions of the eSIP problems in the following examples are computed using the open-source software BET [36].
@@
Figure 6: Heatmaps of solutions of the uniform prior SIP solution for the exponential decay model for $T=2$ computed using tessellation points $\{\lambda_{i}\}_{i=1}^{N_{\jmath}}\in\Lambda$ for $N_{\jmath}\in\{50^{2},100^{2},400^{2},1600^{2}\}$ in order from left to right.
@@
Figure 7: Heatmaps of solutions of the uniform prior SIP solution for the exponential decay model for $T=2$ computed using partitions $\{I_{i}\}_{i=1}^{M_{K}}$ of $\mathcal{D}$ for $M_{K}\in\{12,25,50,100,200\}$ in order from left to right.
The results shown in the lefthand plots in Fig.s 6 and 7 illustrate the analysis that requires $M_{K}$ and $N_{\jmath}$ to become large together in order to improve accuracy.
@@
We solve the eSIP for parameters in a real-world experiment involving falling objects. While the example is simple, it presents challenges involved with formulating and solving an eSIP that generally arise in any application.
@@
There is a closed form solution of the model (4.1), though it is complicated to evaluate. However, using the full model (4.1) has some drawbacks. If we treat $g$, $\rho$, $C_{d}$, $A$, $m$, $H_{0}$, and $V_{0}$ as uncertain parameters, then we are solving an SIP for a distribution on a seven dimensional parameter space from a one dimensional observation. Since the prior disintegrates to a uniform distribution on the six dimensional generalized contours, the eSIP solution will have significant indeterminacy. Alternatively, since $\rho$, $C_{d}$, $m$, and $A$ can be measured independently, we could treat them as random effects. However, if we substitute values from Table 1 into (4.1), we obtain a different model for each type of ball. This means we only have 2-5 samples for the eSIP for each type of ball. Moreover, we would have to deal with the complexity of the constrained eSIP [5].
@@
The number of observations ($N\jmath=17$) is too small to accurately estimate $\hat{p}_{K,M_{K},i}$ directly. We illustrate how to deal with this using simulation-based inference on the distribution of the experimental data by describing two approaches. First, we use a Beta distribution fitted to the observations to generate synthetic data for solving the eSIP. Alternatively, we augment the measured values by adding i.i.d. noise to produce a sufficiently large dataset for the eSIP. This approach amounts to jittering the observations many times to created a smoother empirical distribution function. We use this noisy data to solve the eSIP.
@@
For the first approach, we use maximum likelihood to fit the Beta distribution to the data obtained by shifting the midpoint of $\mathcal{D}$ to $1/2$ then scaling the resulting set to $[0,1]$. The estimated distribution on the unit interval is $\mathrm{Beta}(2.191,2.047)$. We reverse the transformation to obtain a shifted and scaled Beta distribution on $\mathcal{D}$. Finally, $10^{6}$ random samples are generated from this distribution to use for solving the eSIP.
@@
For both approaches, we use $N=27\times 10^{6}$ samples in $\Lambda$ and $M=80$ partitions of $\mathcal{D}$ to compute the solution of the eSIP for the uniform prior. We use a large numbers of samples for the discretization to reduce inaccuracies arising from finite Monte Carlo sampling. We obtain qualitatively similar results using samples on the order of hundreds.
Solving the eSIP.
We show the eSIP solution corresponding to fitting a Beta distribution to the observed data in Fig. 9. The solution locates a region of highest relative probability near the corner $(43\;1\;8.8)^{\top}$, which is largest permissible height and upwards velocity and lowest permissible value for the gravitational constant. Values in this region increase the time of fall, and presumably are compensating for elimination of the effects of drag in the simple model.
Figure 9: Heatmaps of marginal distributions for the solution of the eSIP computed using a parametric fit to the observed data. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$. We indicate the part of $\Lambda$ intersecting the realistic range of uncertainty for $g$ near sea level as a transparent vertical rectangle in the second and third plots.
We show the eSIP solution corresponding to creating “noisy” output data in Fig. 10.
Figure 10: Heatmaps of marginal distributions for the solution of the eSIP computed using noisy data. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$. We indicate the part of $\Lambda$ intersecting the realistic range of uncertainty for $g$ near sea level as a transparent vertical rectangle in the second and third plots.
This solution also places the highest probability near the corner point $(43\;1\;8.8)^{\top}$.
@@
The effective support of the distribution is much smaller than that of the solution computed from fitting the data with a Beta distribution. Further testing shows that decreasing the range of the noise to $q_{i}-.2,q_{i}+.2]$ for each $q_{i}$ produces a solution that very closely resembles the solution obtained with fitted data.
@@
The eSIP solutions in Figures 9 and 10 give cause for concern. Realistically, the variation in the gravitational constant near sea level is on the order of $\pm 0.02\,\text{m/s}^{2}$. Looking at the solution to the eSIP, we see that values within $g\in[9.78,9.82]$ have relatively low probability for either solution.
@@
Fitting a Beta distribution to the data yields $\mathrm{Beta}(1.394,2.030)$ (on the unit interval). We generate $10^{6}$ random samples from this distribution to use for solving the eSIP. We use $N=27\times 10^{6}$ samples in $\Lambda$ and $M=80$ partitions of $\mathcal{D}$ to compute the solution for the uniform prior, see Fig. 11.

### v3

@@ 1.5 Outline of the paper
In Sec. 2, we develop a formulation of nonparametric Bayesian calibration of computer models using disintegration of measures. We describe conditions that guarantee the posterior can be expressed in terms of a conditional density in Sec. 2.4. We describe the maximum entropy property of the posterior corresponding to the uniform prior in Sec. 2.5. In Sec. 2.6, we describe a continuity property of the posterior. In Sec. 3, we describe and analyze an estimator based on reweighting of random samples. In Sec. 4, we present two examples. We develop an alternative accept-reject estimator in Sec. 5. We present a conclusion in Sec. 6. Proofs and additional theoretical developments are provided in Appendix A.
@@ 2 Solution of the Stochastic Inverse Problem
In developing the Bayesian solution of the SIP, we move from the most general formulation of the SIP to a practical formulation by layering the assumptions required for the development. The development is necessarily abstract because the intent is to establish the theoretical framework that provides the basis to compute a robustly accurate empirical estimate of a posterior distribution solving the eSIP that is conditioned on experimental data.
@@ 2.1 Disintegration of measures
Figure 5: Left: we illustrate disintegration when the conditional probabilities are written in terms of densities. The probability on $\Lambda$ is represented by a density shaded in green and the probability on $\mathcal{D}$ is represented by a density shaded in blue. The disintegrated density on $Q^{-1}(q)$ is indicated with a darker shade of green. Right: we illustrate the estimation by random sampling. The observed data is indicated by dark red points on $\mathcal{D}$ and the associated empirical estimator is shown in blue relative to the partition $\{I_{i}\}$. We show the samples $\{\lambda_{i}\}$ in $\Lambda$ as points and $Q^{-1}(I_{i})$ is shaded in gold. Figure is adapted from [47].
@@ 2.2 Bayesian solution of the SIP
To obtain observational data for the eSIP, we compute 200 draws from a Bernoulli random variable with probability $\frac{61}{90}$ of $1$ and use these to compute the estimate $\begin{pmatrix}\hat{P}_{\mathcal{D}}(0)\\ \hat{P}_{\mathcal{D}}(1)\end{pmatrix}=\begin{pmatrix}.34\\ .66\end{pmatrix}$. We compute the posterior corresponding to $\hat{P}_{\mathcal{D}}$. For reasons explained below, in the absence of prior information, we assume a uniform prior,
@@
where $A$ is any set in $\Lambda$ and $\chi_{A}$ is the indicator function for set $A$, as well as the estimated posterior,
@@
While disintegration is a powerful tool [21], its general form can yield extremely “rough” distributions. In order to use disintegration (2.5) to obtain distributions that can be estimated by sampling, we develop the properties of $Q^{-1}$. To this end, we assume the following holds.
@@
The compactness of $\Lambda$ is reasonable because parameters are generally bounded for most process models [^2] We emphasize that $\Lambda$ is not the domain for parameters in a parameterized family of probability distributions.. In fact, determining $\Lambda$ is often an important part of constructing a model, see [47] where the choice of $\Lambda$ for transmission rates for different strains of COVID-19 is discussed. Assuming $\overline{\mathrm{int}\,(\Lambda)}=\Lambda$ and $\mu_{\mathcal{L}}(\partial\Lambda)=0$ are technical assumptions required for estimating the posterior by finite sampling. We illustrate in Fig. 6. The smoothness assumption on $Q$ can be established for broad classes of models since it is typically assumed for computer models. The assumption of GD means that $\Lambda$ can be covered by a finite number of open sets up to a set of measure 0 such that $J_{Q}$ has full rank $m$ on each set. In these sets, no component of $J_{Q}$ can be expressed as a linear combination of its other components. If this is violated, $Q$ should be simplified to obtain a new $Q$ map into a range with smaller dimension.
@@
2.4 Writing the posterior in terms of a conditional density
We develop a formula for the posterior in terms of a conditional density function. We assume,
@@ 2.4 Writing the posterior in terms of a conditional density
Assume Assumptions 2.1, 2.5, 2.7, and 2.8 hold. There is a posterior ${P}_{\Lambda}$ that is absolutely continuous with respect to $\mu_{\Lambda}$ and $dP_{\Lambda}={\rho}_{\Lambda}\,d\mu_{\Lambda}$ with
@@
2.5 The maximum entropy property of the posterior corresponding to a uniform prior
The posterior corresponding to the uniform prior $P_{p}=\frac{1}{\mu_{\Lambda}(\Lambda)}\mu_{\Lambda}$ has density,
@@ 2.5 The maximum entropy property of the posterior corresponding to a uniform prior
We call the posterior corresponding to the uniform prior the posterior conditioned on the uniform prior or the uniform prior posterior.
The uniform prior has a special property that makes it the default choice for solving the SIP in the absence of an informed prior. Namely, it is an expression of the Principle of Insufficient Reason. The translation invariance of the uniform prior respects the fact that since distinct points on the generalized contour $Q^{-1}(q)$ cannot be distinguished using data on $Q$, the probability of points on $Q^{-1}(q)$ should be equal. The next theorem shows that the posterior ${P}_{\Lambda}$ corresponding to the uniform prior is the least biased solution of the SIP in a certain sense. We give the proof in § A.3.
@@
Assume Assumptions 2.1, 2.5, and 2.7 hold. The posterior corresponding to the uniform prior has maximum entropy among all posteriors that are absolutely continuous with respect to $\mu_{\Lambda}$ and whose prior is absolutely continuous with respect to the Lebesgue measure.
@@
In Example 3, the entropy of the posterior for the uniform prior is 2.1746 and the entropy for the posterior for the prior (2.7) is 1.9805.
2.6 Continuity of the posterior
We develop sufficient conditions that guarantee the density of the posterior is continuous a.e. A key ingredient is the a.e. continuity of the surface (Hausdorff) measure of the generalized contours. This depends on the interaction of the generalized contours with the boundary of $\Lambda$. For example, if the boundary has measure larger than 0, there is little chance that the surface measure of the contours is continuous, see Fig. 6.
@@ 2.6 Continuity of the posterior
To analyze the interaction of the generalized contours with the boundary of $\Lambda$, we assume the boundary is a piecewise smooth manifold determined as the level set of a locally smooth function. Specifically, we assume
@@
Assume Assumptions 2.1, 2.5, 2.7, 2.8, and 2.11 hold and the density of the prior $\rho_{\mathrm{p}}$ is continuous a.e. Then, $\tilde{\rho}_{\mathrm{p},\mathcal{D}}(q)$ is continuous in a neighborhood of each point $q\in\mathcal{D}$. In addition, if the density function $\rho_{\mathcal{D}}$ is continuous a.e., then the density ${\rho}_{\Lambda}$ of the posterior $P_{\Lambda}$ corresponding to the prior $P_{\mathrm{p}}$ is continuous in a neighborhood of each point $q\in\mathcal{D}$.
@@
3 Nonparametric estimation of the posterior by reweighting random samples
In this section, we present and analyze a nonparametric estimator of the posterior based on random sampling. The convergence analysis is conducted under the assumptions in Theorem 2.9 and the a.e. continuity of the posterior is crucial.
@@ 3.2 Constructing the estimate
There are three discretizations involved with the eSIP; the number of observed datum $K$, the number of cells $M$ in a partition of $\mathcal{D}$, and the number of cells $N$ in a partition of $\Lambda$. However, limits on the accuracy of empirical density approximations means that $M$ must be related to $K$ in practice. We denote the observed data by $\{q_{i}\}_{i=1}^{\infty}$, which we assume is a random sample generated by a Poisson point process with distribution $P_{\mathcal{D}}$. For $K\geq 1$, we construct an approximation using $\{q_{i}\}_{i=1}^{K}$ that is associated with a partition $\mathcal{I}_{M_{K}}$ of $\mathcal{D}$, where the number of cells $M_{K}$ in $\mathcal{I}_{M_{K}}$ is a function of $K$ that tends monotonically to $\infty$ as $K\to\infty$. We assume the family $\{\mathcal{I}_{M_{K}}\}_{K=1}^{\infty}$ satisfies Assumption 3.1. The approximation is also associated with partitions $\{\mathcal{T}_{N_{\jmath}}\}_{\jmath=1}^{\infty}$ of $\Lambda$, which we assume satisfies Assumption 3.1.
@@
We generalize to other estimators in § 3.5. The approximation for ${P}_{\Lambda}$ is
@@
We write the estimator in a form more conducive to analysis,
@@
An illustration of the estimator is presented in Fig. 5. Beginning with the observed data, we partition the range $\mathcal{D}$ into a collection of cells and use the data to approximate the probability $P_{\mathcal{D}}$ of each cell. The inverse image under $Q$ of each cell in the partition of $\mathcal{D}$ is a “fat” contour. As the cells in the partition of $\mathcal{D}$ become smaller, the fat contours become “thinner”. We compute a sample of points in $\Lambda$ distributed according to the prior. We estimate the probability of the intersection of a fixed set $A$ with a fat contour by taking the ratio of the number of samples in the intersection of a fixed set $A$ with the fat contour and the total number of samples in the fat contour. Finally, we sum up all the estimates over fat contours that intersect $A$, reweighting the estimate in each fat contour with the probability of its corresponding cell in $\mathcal{D}$. The key to convergence of this approach is the a.e. continuity of the surface measure of the contours. When a fat contour is sufficiently thin, the surface measures of almost all the contours in it are approximately the same.
3.3 The basic statistical properties of the estimator
These properties quantify how the estimator converges to the posterior determined by the prior in the limit of large numbers of samples. We use the notation,
@@ 3.3 The basic statistical properties of the estimator
    The estimator is asymptotically unbiased. For $A\in\mathcal{B}_{\Lambda}$ with $\mu_{\Lambda}(\partial A)=0$,
    ${E}\big(\hat{P}_{\Lambda,K,N_{\jmath}}(A)\big)=\sum_{i=1}^{M_{K}}\frac{\breve{p}_{i}^{(K)}}{p_{i}^{(K)}}\big(1-(1-p_{i}^{(K)})^{N_{\jmath}})\,P_{\mathcal{D}}\big(I_{i}^{(K)}\big),$ (3.6)
    and $\lim_{K\to\infty}\lim_{\jmath\to\infty}{{E}\big(\hat{P}_{\Lambda,K,N_{\jmath}}(A)\big))}={P}_{\Lambda}(A)$.
@@
    The estimator is strongly consistent. For $A\in\mathcal{B}_{\Lambda}$ with $\mu_{\Lambda}(\partial A)=0$,
@@
3.4 Asymptotic properties of the error of the estimator
@@
Assume Assumptions 2.1, 2.5, 2.7, 2.8, 2.11, and 3.1 hold and assume the densities $\rho_{\mathrm{p}}$ and $\rho_{\mathcal{D}}$ are continuous a.e. Then, $\lim_{\jmath\to\infty}{E}(T_{3,i,\jmath})=0$ and $\lim_{\jmath\to\infty}\mathrm{Var}\,(T_{3,i,\jmath})=0$.
@@ 3.5 Using other estimators of $\rho_{\mathcal{D}}$
We consider a general estimator $\hat{\rho}_{k,i}=\hat{P}_{\mathcal{D}}\big(I_{i}^{(k)}\big)$ in (3.3), where $\hat{P}_{\mathcal{D}}$ has density $\hat{\rho}_{\mathcal{D}}$ with respect to $\mu_{\mathcal{D}}$. We prove the following theorem in § A.7.
@@
3.6 An important consequence of the analysis
Theorems 3.2, 3.3, 3.4 and 3.5 imply that if sufficient random samples from the prior in $\Lambda$ are used to build the estimator for the posterior, the error of the estimator is determined principally by how well the distribution of the observed data can be estimated. That is the best possible result we can expect. We illustrate this in Example 7.
@@ 4 Examples
We present a couple of simple examples that illustrate the results above. Applications of the eSIP to more complex models with higher dimensional parameter spaces can be found in [10, 12, 14, 13, 53, 36, 15, 16, 18, 54, 6, 43, 47]. The estimates in the following examples are computed using the open-source software BET [37].
@@
Figure 7: Heatmaps of the uniform prior posteriors for the exponential decay model for $T=2$ computed using tessellation points $\{\lambda_{i}\}_{i=1}^{N_{\jmath}}\in\Lambda$ for $N_{\jmath}\in\{50^{2},100^{2},400^{2},1600^{2}\}$ in order from left to right.
@@
Figure 8: Heatmaps of the uniform prior posteriors for the exponential decay model for $T=2$ computed using partitions $\{I_{i}\}_{i=1}^{M_{K}}$ of $\mathcal{D}$ for $M_{K}\in\{12,25,50,100,200\}$ in order from left to right.
The results shown in the left-hand plots in Fig.s 7 and 8 illustrate the analysis that requires $M_{K}$ and $N_{\jmath}$ to become large together in order to improve accuracy.
@@
We calibrate a computer model for falling objects. While the example is simple, it presents challenges involved with formulating and solving an eSIP that generally arise in calibration of computer models.
@@
There is a closed form solution of the model (4.1), though it is complicated to evaluate. However, using the full model (4.1) has some drawbacks. If we treat $g$, $\rho$, $C_{d}$, $A$, $m$, $H_{0}$, and $V_{0}$ as uncertain parameters, then we are estimating a posterior on a seven dimensional parameter space from a one dimensional observation. Since the prior disintegrates to a uniform distribution on the six dimensional generalized contours, the posterior will have significant indeterminacy. Alternatively, since $\rho$, $C_{d}$, $m$, and $A$ can be measured independently, we could treat them as random effects. However, if we substitute values from Table 1 into (4.1), we obtain a different model for each type of ball. This means we only have 2-5 samples for calibration for each type of ball. Moreover, we would have to deal with the complexity of the constrained eSIP [5].
@@
The number of observations ($N\jmath=17$) is too small to accurately estimate $\hat{p}_{K,M_{K},i}$ directly. We illustrate how to deal with this using simulation-based inference on the distribution of the experimental data by describing two approaches. First, we use a Beta distribution fitted to the observations to generate synthetic data for calibration. Alternatively, we augment the measured values by adding i.i.d. noise to produce a sufficiently large dataset for the calibration. This approach amounts to jittering the observations many times to created a smoother empirical distribution function.
@@
For the first approach, we use maximum likelihood to fit the Beta distribution to the data obtained by shifting the midpoint of $\mathcal{D}$ to $1/2$ then scaling the resulting set to $[0,1]$. The estimated distribution on the unit interval is $\mathrm{Beta}(2.191,2.047)$. We reverse the transformation to obtain a shifted and scaled Beta distribution on $\mathcal{D}$. Finally, $10^{6}$ random samples are generated from this distribution to use for the calibration.
@@
For both approaches, we use $N=27\times 10^{6}$ samples in $\Lambda$ and $M=80$ partitions of $\mathcal{D}$ to compute the posterior for the uniform prior. We use a large numbers of samples for the estimator to reduce inaccuracies arising from finite Monte Carlo sampling, following Theorems 3.2, 3.3, 3.4 and 3.5. We obtain qualitatively similar results using samples on the order of hundreds.
Calibrating the computer model using all data.
We show the posterior corresponding to fitting a Beta distribution to the observed data in Fig. 10. The posterior locates a region of highest relative probability near the corner $(43\;1\;8.8)^{\top}$, which is largest permissible height and upwards velocity and lowest permissible value for the gravitational constant. Values in this region increase the time of fall, and presumably are compensating for elimination of the effects of drag in the simple model.
Figure 10: Heatmaps of marginal distributions for the posterior computed using a parametric fit to the observed data. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$. We indicate the part of $\Lambda$ intersecting the realistic range of uncertainty for $g$ near sea level as a transparent vertical rectangle in the second and third plots.
We show the posterior corresponding to creating “noisy” output data in Fig. 11.
Figure 11: Heatmaps of marginal distributions for the posterior computed using noisy data. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$. We indicate the part of $\Lambda$ intersecting the realistic range of uncertainty for $g$ near sea level as a transparent vertical rectangle in the second and third plots.
This posterior also places the highest probability near the corner point $(43\;1\;8.8)^{\top}$.
@@
The effective support of this posterior is much smaller than that of the posterior computed from fitting the data with a Beta distribution. Further testing shows that decreasing the range of the noise to $q_{i}-.2,q_{i}+.2]$ for each $q_{i}$ produces a posterior that very closely resembles the posterior obtained with fitted data.
@@
The posteriors in Figures 10 and 11 give cause for concern. Realistically, the variation in the gravitational constant near sea level is on the order of $\pm 0.02\,\text{m/s}^{2}$. Either posterior implies that values within $g\in[9.78,9.82]$ have relatively low probability.
@@
Fitting a Beta distribution to the data yields $\mathrm{Beta}(1.394,2.030)$ (on the unit interval). We generate $10^{6}$ random samples from this distribution to use for solving the eSIP. We use $N=27\times 10^{6}$ samples in $\Lambda$ and $M=80$ partitions of $\mathcal{D}$ to compute uniform prior posterior, see Fig. 12.